# MFC 5.0: An exascale many-physics flow solver

### v3

@@ 4.3.3 Exascale capabilities and beyond
The communication patterns in MFC also result in high strong scaling efficiencies. Figure 7 shows the strong scaling performance of MFC on OLCF Summit and OLCF Frontier and compares weak scaling with and without GPU-aware MPI. The problem sizes in fig. 7 (a) correspond to approximately 100% and 50% usage of available GPU memory per device in the base case. In the base case, with 8M grid cells per device, strong scaling efficiencies of 79% and 51% are observed on OLCF Summit when increasing the device count by a factor of 8 or 16 without GPU-aware MPI. For the case with 32 million grid cells per device in the base case, strong scaling efficiencies of 81% and 77% when increasing the device count by a factor of 8 and 16 without GPU-aware MPI. The higher efficiencies seen with the AMD MI250X on OLCF Frontier are partly due to the larger problem size’s increased computation-to-communication ratio and the improved internode communication on Frontier. With GPU-aware MPI, the strong scaling efficiencies on OLCF Summit increase to 84% and 60% when increasing the device count by a factor of 8 and 16. On Frontier, the efficiencies are 96% and 92% when the device count is increased by the same factors.
@@ 4.3.4 Parallel I/O at extreme scales
Parallel I/O is implemented using the MPI I/O library to facilitate shared file parallel I/O and file-per-process parallel I/O. Shared file parallel I/O was sufficient for handling up to 10K processes, but a significant slowdown in I/O times was observed when scaling to over 65K processes on OLCF Frontier. For simulations that use over 10K processes, each process writes a file for each I/O operation. To alleviate the file system pressure resulting from metadata creation for tens of thousands of files, MFC accesses the file system in waves of 128 processes separated by a fixed number of floating point operations.
@@ 4.3.5 Software resilience and continuous integration
Each pull request to MFC goes through a suite of continuous integration actions. The correctness, performance, and code quality are evaluated through this continuous integration suite. MFC’s test suite of over 300 cases is compiled and run using GNU and Intel compilers on MacOS and Ubuntu operating systems on CPUs on GitHub runners via GitHub actions. The same tests are run using NVHPC on CPUs and NVIDIA GPUs and with CCE on AMD GPUs via self-hosted runners. All of these tests are compared to the same set of golden files to ensure that the code in the main branch of MFC provides the same results on CPUs and GPUs, regardless of compiler or operating system. The coverage of the test suite is evaluated using Codecov to help ensure coverage of the test suite for additional (or removed) code. Each pull request is benchmarked against the MFC master branch on CPUs and GPUs to prevent performance regressions; it also checks for compiler warnings and formatting to ensure code consistency.
@@ 4.3.6 Metaprogramming
Metaprogramming is implemented using the Fortran preprocessor Fypp [84]. Fypp is a Python-based preprocessor that is used to inline subroutines via programmer directives and collapse repetitive code. MFC uses Fypp for several tasks, including automating deep copies of derived types for allocation on the GPU device. NVIDIA’s NVHPC Fortran compiler automatically allocates derived types on the device, but CCE adheres closer to the OpenACC technical specification (3.2 at the time of writing) and requires manual allocation. If one holds many host variables, this process is cumbersome. Instead, we use the code in LABEL:lst:deepCopy to inline the source code via @:ACC_SETUP_VFs(var1,var2,…,varN) after allocating the vector fields var1,var2,…,varN on the host. This macro performs deep copies, reducing the number of lines required to initialize the variables at runtime.
Listing 2: A Fortran+Fypp macro that manually deep copies derived types from the host to the device.
#:def acc_setup_vfs(\*args)
@@
!\$acc enter data copyin(\${arg}\$)
!\$acc enter data copyin(\${arg}\$%vf)
@@
!\$acc enter data copyin(\${arg}\$%vf(macros_setup_vfs_i))
!\$acc enter data create(\${arg}\$%vf(macros_setup_vfs_i)%sf)
@@
Fypp simplifies module initialization by wrapping allocate and deallocate into macros that allocate the variable and perform the enter data create required when using OpenACC for GPU offloading. LABEL:lst:memManag shows an example macro. Variables are allocated and deallocated using @:ALLOCATE(var1,var2,…,varN) and @:DEALLOCATE(var1,var2,…,varN). The code length is further reduced by using Fypp to abstract away repetitive code via preprocessor loops [85].
Listing 3: An Fypp macro for allocating and initializing a variable with OpenACC.
@@
allocate (\${’, .join(args)}\$)
!\$acc enter data create(\${’, ’.join(args)}\$)
@@
deallocate (\${’, ’.join(args)}\$)
!\$acc exit data delete(\${’, ’.join(args)}\$)
@@
Metaprogramming sets problem parameters as constant parameters in case files. This option is called case optimization, enabled via the flag --case-optimization at run time. Case optimization reduces register pressure by providing the compiler with the parameters of a configuration. On CPUs, specifying the problem parameters at compile-time results in a two-times speedup of the code. On GPUs, case optimization decreases run time by about a factor of ten. Variables are precriped in a file with the Fypp code #:set var = val. This file is included in the Fortran source code; compile-time-known parameters are declared using the Fypp logic of LABEL:lst:fyppDec.
Listing 4: Fortran+Fypp code the declaration of configuration-specific parameters for compile-time optimization.
@@ 4.3.7 Code generation for efficient reactions and thermodynamics
MFC now simulates chemically reacting flows. We accommodate this feature via combustion routines that evaluate the chemical source terms and species thermodynamics. Comprehensive libraries, such as Cantera, provide such routines. However, these exist in a different compilation unit to the flow solver, so they cannot easily be offloaded via the OpenACC directive-based strategy of section 4.3.1. We address this challenge through code generation, producing a computational representation of thermochemistry at the same abstraction level as MFC’s compressible flow solver. This strategy obviates the need to optimize comprehensive combustion libraries at link time. MFC uses Pyrometheus [58, 86] for code generation, generating thermochemistry code in OpenACC-decorated Fortran.
Pyrometheus is based on a symbolic representation of the combustion formulation that takes mechanism parameters from Cantera. The symbolic representation is mapped to a user-specified target language. For Fortran, the generated code is a module that can be easily integrated and offloaded with MFC. The mapping process decorates the routines with OpenACC statements for GPU offloading, so the generated code does not need to be modified a posterior. The generated code unrolls inner loops over species and reactions and prepares compile-time constants such as Arrhenius parameters. This optimization meaningfully reduces kernel runtime [86]. In a similar vein to case optimization of section 4.3.6, Pyrometheus assigns compile-time-known bounds to the arrays in the generated code, enabling optimization of register allocations and decreasing runtime by about a factor of 5 for GPU kernels.
Pyrometheus is verified against Cantera, though it is itself thoroughly tested. The testing suite is language-agnostic, so the generated Fortran code is as accurate as a C-based alternative. The generated code is mechanism-specific, but because code generation is free in comparison with simulation time, multiple mechanisms can be preprocessed. For MFC, we generate code for established combustion mechanisms for hydrogen [63] and methane combustion [87], though we are not limited to these.
@@ 5.1 Shock–bubble-cloud interaction
[…]
@@
[…]
@@ 5.2 Shock–bubble interaction
[…]
@@ 5.3 Taylor–Green vortex
[…]
@@ 5.4 Supersonic gas injection and plume dynamics
Three jets inject into the domain via Dirichlet boundary conditions. The domain has size $[0,16D]\times[0,10D]\times[0,10D]$ where $D$ is the jet diameter and the first dimension is the mean flow direction (horizontal in fig. 11). The domain is discretized by $600\times 400\times 400$ grid cells. The jets have incoming velocity ($V_{\mathrm{jet}}$), pressure, and density that satisfy the Mach 1.5 normal shock conditions; they are organized in an equilateral triangle. The jet Reynolds number is $\mbox{\text{Re}}=\rho V_{\mathrm{jet}}D/\mu=2.5\times 10^{6}$.
@@
The simulation is executed until dimensionless time $40D/(\mbox{\text{Ma}}\,c)$, where Ma is the Mach number and $c$ is the ambient speed of sound. The last time step is visualized in fig. 11. The visualization shows a volume rendering of the injected fluid after it has meaningfully intruded the domain.
@@ 6 Solvers with some shared capabilities
Other open-source codes for CFD exist. These and their relative trade-offs compared to MFC 5.0 are briefly discussed here. URANOS-2.0 (De Vanna and Baldan [89]) is a modern Fortran code that uses OpenACC for GPU offloading to NVIDIA and AMD GPUs. URANOS-2.0 contains some of MFC’s shock-capturing capabilities; it also has compressible turbulence models. STREAmS-2 (Sathyanarayana et al. [90]) is also a Fortran code for compressible flow and uses CUDA Fortran and hipFORT to offload to NVIDIA and AMD GPUs. There is also an OpenMP port of STREAmS-2 that makes it portable to Intel GPUs. AFiD-GPU of Zhu et al. [91] is similar in this context. Neko (Jansson et al. [92]) is a particularly modern object-oriented solver that uses multiple levels of abstraction for offloading to GPUs as well as more extensive extensions for vector processors and FPGAs. The above solvers have offloading capabilities similar to those of MFC, though the physical models and their sophistication differ. MFC offers a broad range of modeling capabilities and numerics for the broad use of multi-phase and -species flows,
@@ 7 Conclusion
Since MFC 3.0 [1], the code has transitioned from a specialized research code to an exascale-capable framework for multiphysics and multiphase flows. The implementation of physical models, numerical methods, software infrastructure, and high-performance computing tools embodies this transition.
@@
Numerical advancements accompany MFC’s extended feature set. TENO and WENO-Z shock-capturing schemes meaningfully reduce dispersion error compared to conventional WENO schemes. The Strang-split particle solver maintains second-order temporal accuracy for stiff ODE systems. The implementation of a low-Mach preconditioner reduces numerical dissipation for Mach numbers below $0.1$. A ghost-cell immersed boundary method is now available. This method supports complex 2D and 3D geometries that can be imported as level sets or STL files.

### v4

@@ 4.3.3 Exascale capabilities and beyond
The communication patterns in MFC lead to high strong scaling efficiencies. Specifically, the communication involves only a small halo region around each uniformly sized domain decomposition chunk, and the relatively large amount of computation associated with each grid point. The decomposed grid chunks, assigned one-to-one per MPI rank, are not expanded or contracted. Since each computing device handles the same number of elements, the halo exchange cost is the same for each MPI rank, which also explains the ideal weak scaling behavior: The number of domain decomposition chunks increases linearly with problem size, so there is no opportunity for weak scaling inefficiencies.
Figure 18 shows the strong scaling performance of MFC on OLCF Summit and OLCF Frontier and compares weak scaling with and without GPU-aware MPI. The problem sizes in fig. 18 (a) correspond to approximately 100% and 50% usage of available GPU memory per device in the base case. In the base case, with 8M grid cells per device, strong scaling efficiencies of 79% and 51% are observed on OLCF Summit when increasing the device count by a factor of 8 or 16 without GPU-aware MPI. For the case with 32 million grid cells per device in the base case, strong scaling efficiencies of 81% and 77% when increasing the device count by a factor of 8 and 16 without GPU-aware MPI. The higher efficiencies seen with the AMD MI250X on OLCF Frontier are partly due to the increased computation-to-communication ratio resulting from larger problem sizes and the improved internode communication on Frontier. With GPU-aware MPI, the strong scaling efficiencies on OLCF Summit increase to 84% and 60% when increasing the device count by a factor of 8 and 16. On Frontier, the efficiencies are 96% and 92% when the device count is increased by the same factor.
@@ 4.3.4 Parallel I/O at extreme scales
Parallel I/O is implemented using the MPI I/O library to facilitate shared file parallel I/O and file-per-process parallel I/O. Shared file parallel I/O was sufficient for handling up to 10K processes, but a significant slowdown in I/O times was observed when scaling to over 65K processes on OLCF Frontier. For simulations that use over 10K processes, each process writes a file for each I/O operation. To alleviate file system pressure resulting from metadata creation for tens of thousands of files, MFC accesses the file system in waves of 128 processes, separated by a fixed number of floating-point operations.
@@ 4.3.5 Software resilience and continuous integration
Each pull request to MFC goes through a suite of continuous integration actions. The correctness, performance, and code quality are evaluated through this continuous integration suite. MFC’s test suite, comprising over 500 cases, is compiled and run using GNU and Intel compilers on macOS and Ubuntu operating systems, utilizing GitHub Actions and Runners. The same tests are run using NVHPC on CPUs and NVIDIA GPUs, as well as with CCE on AMD GPUs via self-hosted runners. All of these tests are compared to the same set of golden files to ensure that MFC’s code provides the same results on CPUs and GPUs, regardless of compiler or operating system. The coverage of the test suite is evaluated using Codecov to help ensure coverage of the test suite for additional (or removed) code. Each pull request is benchmarked against the main MFC branch on CPUs and GPUs to prevent performance regressions; it also checks for compiler warnings and formatting to ensure code consistency.
@@ 4.3.6 Metaprogramming
Metaprogramming is implemented using the Fortran preprocessor Fypp [fypp]. Fypp is a Python-based preprocessor that inlines subroutines via directives and collapses repetitive code. MFC uses Fypp for several tasks, including automating deep copies of derived types for allocation on the GPU device. NVIDIA’s NVHPC Fortran compiler automatically allocates derived types on the device, whereas CCE adheres more closely to the OpenACC technical specification (version 3.2 at the time of writing) and requires manual allocation. If one holds many host variables, this process is cumbersome.
LABEL:lst:deepCopy shows how Fypp is used to generate repetitive directives required for deep copying derived types. Without metaprogramming, each component of a derived type would need to be copied individually. The macro @:SETUP_VFs(var1,var2,…,varN) is used after allocating the vector fields var1,var2,…,varN on the host. This macro automatically traverses the derived type fields and inserts the correct enter data and create statements, ensuring that all vector fields and their subfields are allocated on the device with minimal manual coding. This macro also performs deep copies, reducing the number of lines required to initialize the variables at runtime.
Listing 3: A Fortran+Fypp macro that manually deep copies derived types from the host to the device. In the loop, Fypp expands @:SETUP_VFs(var1, var2, …) into repeated OpenACC or OpenMP directives. \$:GPU_ENTER_DATA(copyin=’[]’ reduces to !\$acc enter data copyin or !\$omp target enter data copyin depending on the desired GPU offload too. For each variable, the macro first copies the variable and its vector field array to the device. If the array is allocated, the loop iterates over its bounds, and if any scalar field pointers are associated, it performs further deep copies. This ensures a complete mirror of the host data structure on the GPU.
#:def SETUP_VFs(*args)
@@
\$:GPU_ENTER_DATA(copyin=(’[’ + arg + ’]’))
\$:GPU_ENTER_DATA(copyin=(’[’ + arg + ’%vf]’))
@@
\$:GPU_ENTER_DATA(copyin=(’[’ + arg + ’%vf(macros_setup_vfs_i)]’))
\$:GPU_ENTER_DATA(copyin=(’[’ + arg + ’%vf(macros_setup_vfs_i)%sf]’))
@@
Fypp simplifies memory management by wrapping allocate and deallocate into macros that allocate the variable and perform the necessary data movement required when using OpenACC or OpenMP for GPU offloading. LABEL:lst:memManag shows this abstraction. A call such as @:ALLOCATE(u, v, w) expands to allocate the three variables and register them on the GPU device. The code length is further reduced by using Fypp to abstract away repetitive code via preprocessor loops [radhakrishnan24].
Listing 4: Fypp macros for automatically allocating and deallocating memory with the associated directives for GPU memory management. \$:GPU_ENTER_DATA(create=’[]’) reduces to !\$acc enter data create or !\$omp target enter data create and \$:GPU_EXIT_DATA(delete=’[]’) reduces to !\$acc exit data delete or !\$omp target exit data delete depending on the desired GPU offload tool.
@@
#:set allocated_variables = ’, ’.join(args)
allocate (\${allocated_variables}\$)
#:set cleaned = strip_parenthesis(allocated_variables)
#:set joined = ’, ’.join(cleaned)
\$:GPU_ENTER_DATA(create=’[’ + joined + ’]’)
@@
#:set allocated_variables = ’, ’.join(args)
\$:GPU_EXIT_DATA(delete=(’[’ + allocated_variables + ’]’))
deallocate (\${allocated_variables}\$)
@@
Metaprogramming sets problem parameters as constant parameters in case files. This option is called case optimization, enabled via the flag --case-optimization at run time. Case optimization reduces register pressure by providing the compiler with the parameters of a configuration. On CPUs, specifying the problem parameters at compile time results in a two-fold speedup of the code. On GPUs, case optimization decreases run time by about a factor of ten.
LABEL:lst:fyppDec demonstrates case optimization: when the option is enabled, configuration parameters are converted into compile-time constants (parameter). Otherwise, they remain runtime variables. This small change enables the compiler to perform constant folding and loop unrolling, thereby reducing register pressure and improving performance. Specifically, variables are prescribed in a file with the Fypp code #:set var = val. This file is included in the Fortran source code; compile-time-known parameters are declared using the Fypp logic of LABEL:lst:fyppDec.
Listing 5: Fortran+Fypp code the declaration of configuration-specific parameters for compile-time optimization.
@@
In summary, Fypp-style metaprogramming enables MFC to automate deep copies of complex data structures, simplify memory management with consistent macros, and facilitate compile-time specialization of problem parameters. These techniques reduce boilerplate code, improve maintainability, and deliver performance improvement on CPUs and GPU devices.
@@ 4.3.7 Code generation for efficient reactions and thermodynamics
MFC now simulates chemically reacting flows. We accommodate this feature via combustion routines that evaluate the chemical source terms and species thermodynamics. Comprehensive libraries, such as Cantera, provide such routines. However, these exist in a different compilation unit from the flow solver, so they cannot easily be offloaded via the OpenACC directive-based strategy of section 4.3.1. We address this challenge through code generation, producing a computational representation of thermochemistry at the same level of abstraction as MFC’s compressible flow solver. This strategy obviates the need to optimize comprehensive combustion libraries at link time. MFC uses Pyrometheus [ref:cisneros2022, cisneros25] for code generation, generating thermochemistry code in OpenACC-decorated Fortran.
Pyrometheus is based on a symbolic representation of the combustion formulation that takes mechanism parameters from Cantera. The symbolic representation is mapped to a user-specified target language. For Fortran, the generated code is a module that can be easily integrated and offloaded with MFC. The mapping process decorates the routines with OpenACC statements for GPU offloading, ensuring that the generated code does not require post-hoc modifications. The generated code unrolls inner loops over species and reactions and prepares compile-time constants such as Arrhenius parameters. This optimization meaningfully reduces kernel runtime [cisneros25]. In a similar vein to case optimization of section 4.3.6, Pyrometheus assigns compile-time-known bounds to the arrays in the generated code, enabling the optimization of register allocations and reducing runtime by approximately a factor of 5 for GPU kernels.
Pyrometheus is verified against Cantera, although it is thoroughly tested itself. The testing suite is language-agnostic, so the generated Fortran code is as accurate as a C-based alternative. The generated code is mechanism-specific, but because code generation is faster than simulation time, multiple mechanisms can be preprocessed. For MFC, we generate code for established combustion mechanisms for hydrogen [ref:Saxena2006] and methane combustion [ref:gri30], though we are not limited to these.
@@ 5 Example simulations
Example simulations using MFC 5.0 are included below. Many of these could not be conducted using prior MFC releases, or would have been prohibitively expansive, as MFC 5.0 has improved distributed computation and compiler and non-NVIDIA GPU hardware support has been added with MFC 5.0.
@@ 5.1 Shock–bubble-cloud interaction
The first example simulation shows the interaction between a cloud of $75$ randomly placed $3\text{\,}\mathrm{mm}\text{/}$ bubbles and a shockwave in water. The computational domain is $9D,9D]\times[0,10D]\times[0,10D]$ before grid stretching and is discretized as $(N_{x},N_{y},N_{z})=(2000,1000,1000)$, or about 2B grid points. The grid points are stretched away from the bubbles to avoid boundary effects via
@@
where $a_{x}$ is the stretching parameter, $L$ is the domain length, and $x_{a}$ and $x_{b}$ control the stretching location.
Figure 19 shows the $\alpha=0.5$ isosurface at increasing points in time from left to right. The detail view shows the grid resolution around one of the collapsing bubbles. The initial condition has 100 uniform-sized grid points across each bubble diameter. The simulation was performed using 1024 AMD MI250X GCDs (512 AMD MI250X GPUs) on OLCF Frontier and completed in approximately $30\text{\,}\mathrm{min}\text{/}$.
@@ 5.2 Shock–bubble interaction
The second example simulation shows the interaction between a helium bubble in air impinged by a shock wave. The computational domain has spatial extents $0,20D]\times[-5D,5D]\times[-5D,5D]$ before grid stretching and is discretized with $(N_{x},N_{y},N_{z})=(3200,1600,1600)$, which is over 8B grid points. The domain boundaries are moved far away from the bubble to avoid boundary effects using grid stretching (described in section 5.1). Figure 20 shows the $\alpha=0.5$ isosurface, which represents the interface between air and helium, at dimensionless times $t^{*}=tU/D=1.5$, $3.25$, and $5.0$ for shock speed $U$ and bubble diameter $D$. The isosurface is colored by velocity magnitude, with darker colors corresponding to higher velocities. The bubble is resolved with 640 grid cells in its initial diameter. The simulation was performed using 144 NVIDIA H200 GPUs in $16\text{\,}\mathrm{h}\text{/}$.
@@ 5.3 Taylor–Green vortex
The third example simulation is a $\mbox{\text{Re}}=1600$, $\mbox{\text{Ma}}=0.1$ Taylor–Green vortex. The initial condition follows that of koen2013. The computational domain is a cube with side lengths of $L=2\pi$ and is discretized with $N_{x}=N_{y}=N_{z}=1600$, resulting in approximately 4 billion grid points. Figure 21 shows the isosurface with zero Q-criterion colored by vorticity magnitude at dimensionless times $t/t_{c}=4$, $12$, and $20,$ with darker colors corresponding to higher vorticity magnitudes. The convective timescale $t_{c}$ is defined as $t_{c}=L/\left(c_{0}\mbox{\text{Ma}}\right),$ where $c_{0}$ is the free stream speed of sound. The simulation was performed using 144 NVIDIA H200 GPUs in $28\text{\,}\mathrm{h}\text{/}$.
@@ 5.4 Supersonic gas injection and plume dynamics
Three jets inject into the domain via Dirichlet boundary conditions. The domain has size $0,16D]\times[0,10D]\times[0,10D]$ where $D$ is the jet diameter and the first dimension is the mean flow direction (horizontal in fig. 22). The domain is discretized into a $600\times 400\times 400$ grid of cells. The jets have an incoming velocity ($V_{\mathrm{jet}}$), pressure, and density that satisfy the Mach 1.5 normal shock conditions; they are organized in an equilateral triangle. The jet Reynolds number is $\mbox{\text{Re}}=\rho V_{\mathrm{jet}}D/\mu=2.5\times 10^{6}$.
@@
The simulation is executed until dimensionless time $40D/(\mbox{\text{Ma}}\,c)$, where Ma is the Mach number and $c$ is the ambient speed of sound. The last time step is visualized in fig. 22. The visualization shows a volume rendering of the injected fluid after it has meaningfully intruded into the domain.
5.5 Validation and verification
The presented examples above are for demonstration of simulation capability and integrity. The shock–bubble dynamics cases of sections 5.1 and 5.2, including the model and numerics, were validated at the scale of single-bubble dynamics by bryngelson21 and the cited art within. The numerics themselves have been validated in prior works, such as johnsen2006implementation. The Taylor–Green vortex of section 5.3 is validated against standard turbulence statistics. The jet case in section 5.4 was verified through a grid study, but is primarily presented as an example of computational readiness.
@@ 6 Solvers with some shared capabilities
Other open-source codes for CFD exist. These and their relative trade-offs compared to MFC 5.0 are briefly discussed here. URANOS-2.0 (devanna24) is a modern Fortran code that uses OpenACC for GPU offloading to NVIDIA and AMD GPUs. URANOS-2.0 incorporates some of MFC’s shock-capturing capabilities and features compressible turbulence models. STREAmS-2 (sathyanarayana2023) is also a Fortran code for compressible flow, utilizing CUDA Fortran and hipFORT to offload computations to NVIDIA and AMD GPUs. There is also an OpenMP port of STREAmS-2 that makes it portable to Intel GPUs. AFiD-GPU of zhu2018 is similar in this context. Neko (jansson2024) is a particularly modern object-oriented solver that uses multiple levels of abstraction for offloading to GPUs as well as more extensive extensions for vector processors and FPGAs. The above solvers have offloading capabilities similar to those of MFC, although their physical models and sophistication differ. MFC offers a broad range of modeling capabilities and numerical methods for simulating multi-phase and multi-species flows.
@@ 7 Conclusion
Since MFC 3.0 [bryngelson21], the code has transitioned from a specialized research code to an exascale-capable framework for multiphysics and multiphase flows. The implementation of physical models, numerical methods, software infrastructure, and high-performance computing tools embodies this transition.
@@
Numerical advancements accompany MFC’s extended feature set. For shock-capturing, the implementation of TENO, WENO-Z, and higher-order reconstructions provides methods that are well-established to reduce the numerical dissipation errors inherent in conventional, lower-order schemes [shu1997essentially, leveque2002finite, borges2008improved, fu2016family]. The Strang-split particle solver with the adaptive time stepping algorithm was implemented to improve numerical efficiency for stiff sub-grid dynamics while ensuring second-order temporal accuracy [strang1968construction, hairer93solving]. The known inaccuracy of Riemann solvers in the low-Mach number limit [guillard1999behaviour] is addressed with two implemented treatments that reduce numerical dissipation, following the approaches of Thornber et al. [thornber2008improved] and Chen et al. [chen2022anti]. A ghost-cell immersed boundary method is now available. This method supports complex 2D and 3D geometries that can be imported as level sets or STL files.